OpenStudy (anonymous):

Help with simplifying! (5−√(10−5x)) * (2+√(10−5x))

OpenStudy (solomonzelman):

OK do you know this? \[(a+b)(a-b)=a^2-b^2\]and the middle term cancels

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Do you see the same thing in your problem?

OpenStudy (anonymous):

=/ the radicals make it confusing for me

OpenStudy (solomonzelman):

\[(5−\sqrt{10-5x}) \times (2+\sqrt{10-5x})\] Sorry it won't cancel, just expand the parenthesis.

OpenStudy (anonymous):

so